# M4-SAR: A Multi-Resolution, Multi-Polarization, Multi-Scene, Multi-Source Dataset and Benchmark for optical-SAR Fusion Object Detection

### v2

@@ III-C Data Processing
We combined the red, green, and blue channels to generate standard optical images using data processing software. The optical images were provided in three spatial resolutions: 10m, 20m, and 60m. To highlight the complementary characteristics of optical and SAR images, we retained only the high-resolution (10m) and low-resolution (60m) variants. For the SAR data, we first applied multi-look averaging to the SLC images in both the azimuth and range directions to reduce speckle effects and improve radiometric resolution. Subsequently, filtering techniques were employed to suppress speckle noise. Using the corresponding DEM data, we geocoded and radiometrically calibrated the SAR images and conducted geometric alignment and correction to improve spatial localization precision. After preprocessing, the optical and SAR images were aligned based on their geographic coordinates. As the overlap between the two modalities was partial, the aligned image pairs were cropped into $\times1,024$ patches, and only those with valid overlapping regions were retained for subsequent annotation. Notably, this initial cropping was intended to facilitate efficient manual annotation. After annotation was completed, we further partitioned the annotated images into non-overlapping $\times512$ tiles to meet the demands of different downstream applications.
@@
III-D Data Annotation and Label Mapping
As the acquisition times of Sentinel-1 [44] and Sentinel-2 [7] are not perfectly synchronized, instance-level alignment of dynamic targets (e.g., ships) becomes difficult, which compromises the accuracy of label transfer. To address this issue, we focus on static and semantically meaningful targets, such as bridges, harbors, oil tanks, playgrounds, airports, and wind turbines, and annotate them using oriented bounding boxes. SAR’s unique imaging characteristics cause ambiguous target boundaries, challenging accurate annotation by non-experts. In contrast, optical images (cloudless) provide clearer visual cues that facilitate both target identification and annotation transfer. With precise alignment, annotations derived from optical images can be reliably transferred to SAR images, thereby significantly reducing manual annotation costs.
Given the scale of the dataset, fully manual labeling is both time-consuming and labor-intensive. To address this challenge, we propose a semi-supervised annotation strategy, as illustrated in Fig. 2 (c). The procedure involves the following steps: First, 5% of the entire cloudless optical images are manually annotated, ensuring that each selected image contains all six target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. This subset establishes a high-quality foundation for subsequent annotation. Once manually labeled, a detector [16] is trained using this subset (with the same data used for both training and evaluation during the initial phase). The trained model is then employed to generate pseudo-labels for an additional 5% of the data, which are subsequently refined through manual correction and added to the labeled set. This updated set is used to retrain the detector in the next iteration. The annotation process is repeated iteratively until the full dataset is labeled. During this procedure, images without valid target instances are excluded to enhance annotation efficiency of the dataset.
During the annotation process, we use cloud-free images to ensure annotation accuracy. After annotation is complete, the cloud-free images are replaced with the corresponding cloud-containing images. Based on instance-level alignment, we achieve high-quality label transfer, enabling accurate projection of annotation information from optical images to their corresponding SAR counterparts. As a result, we construct M4-SAR, the large-scale dataset for optical-SAR fusion object detection, which offers a solid foundation for subsequent method evaluation and algorithm development.
III-E Statistical Analysis of Different Modalities
As shown in Table I, our proposed M4-SAR dataset comprises 112,174 image pairs, 981,862 instances, and six representative target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. Each image has a resolution of $\times512$ pixels and contains an average of 8.75 instances, highlighting the dense object distribution typical in remote sensing scenes. As illustrated in Fig. 1 (c), rotated bounding boxes offer significantly higher localization accuracy compared to horizontal bounding boxes.
Statistical analysis of category-specific attributes in the M4-SAR dataset is presented as follows:
@@ III-E Statistical Analysis of Different Modalities
  • Aspect Ratio: The aspect ratio of an object plays a critical role in detection accuracy. Fig. 1 (b) illustrates the distribution of aspect ratios across categories. When the aspect ratio approaches 1, targets tend to appear square, potentially leading to recognition ambiguities. Conversely, extremely large aspect ratios increase the model’s sensitivity to object shapes, thus posing additional challenges to accurate detection.
@@
  • Angle: All target orientation angles are normalized to $\theta$ ($\theta\in[0,\pi/2)$), and their distributions are visualized in Fig. 3 (a). Notably, offshore wind turbines tend to exhibit square-like appearances, with orientation angles predominantly distributed around 0 and 90 due to the absence of surrounding reference structures and the top-down perspective of satellite imagery. Although oil tanks are circular, their labeled angles are contextually defined based on surrounding spatial features. This highlights the importance of contextual cues in remote sensing images and reflects the task’s sensitivity and complexity.
@@
  • Number of Instances: Fig. 3 (b) displays the distribution of instances by category, revealing that bridges and oil tanks constitute the largest proportions in the dataset. This imbalance reflects the actual occurrence frequency of typical targets in coastal regions: in such areas, static infrastructure such as bridges and oil tanks dominates satellite imagery. The relatively low number of port instances is primarily due to their extensive spatial coverage, with a single port typically corresponding to a single annotated instance.
  • Average Area of Objects: Fig. 3 (c) shows the average pixel area per category, revealing the scale variability among different object categories in the dataset.
@@
Figure 4: (a) optical-SAR domain gap. (b) Overall framework of the proposed end-to-end optical-SAR fusion object detection (E2E-OSDet). Architectural details of the proposed filter augment module (c), cross-modal Mamba interaction module (d), and area-attention fusion module (e).
IV Method
IV-A optical-SAR Fusion Detection Tools
@@ IV Method
Due to the lack of publicly available methods specifically designed for optical-SAR fusion object detection, conducting a fair performance comparison remains challenging. To address this limitation, we propose a Multi-Source Rotated Detector (MSRODet) built upon existing multi-source detection frameworks (e.g., RGB-infrared, RGB-depth). MSRODet integrates several representative methods, including CFT [36], CLANet [14], CSSA [2], CMADet [39], ICAFusion [38], and MMIDet [60], thereby providing a standardized and accessible benchmark for evaluating optical-SAR fusion object detection approaches. As shown in Fig. 4 (a), empirical analysis reveals that significant domain gaps remain a primary obstacle to effective optical-SAR fusion. These discrepancies primarily arise from fundamental differences in visual characteristics and imaging mechanisms between optical and SAR modalities. Consequently, designing detection frameworks specifically tailored to optical-SAR fusion is essential for enhancing overall detection performance.
IV-B End-to-End optical-SAR Fusion Object Detection
IV-B1 Overall Framework
@@
Multi-source fusion for object detection remains relatively underexplored in the context of optical-SAR integration, and notable challenges persist due to the significant domain discrepancies between the two modalities. As illustrated in Fig. 4 (b), we propose an end-to-end optical-SAR fusion framework, E2E-OSDet, which explicitly tackles these cross-modal challenges at three complementary levels: data input, domain alignment, and feature fusion.
IV-B2 Filter Augment Module
To address the inherent domain disparities between optical and SAR images, caused by differences in imaging mechanisms and feature distributions, we propose the Filter Augment Module (FAM), a novel preprocessing framework designed to harmonize multi-source inputs for robust object detection. FAM incorporates a suite of classical image filtering operators, including HOG [4], Canny [1], Haar [45], Grad [21], and WST [34]. These filters extract complementary low-level cues, such as edges, gradients, and textures, under diverse imaging conditions. By projecting the sparse, low-dimensional features of both optical and SAR images into a high-dimensional, discriminative feature space, FAM reduces cross-modal domain gaps and enhances the alignment of modality-specific characteristics. Specifically, given an optical image $I_{O}\in\mathbb{R}^{H\times W\times 3}$ and a SAR image $I_{S}\in\mathbb{R}^{H\times W\times 3}$, FAM generates a set of enhanced representations $F_{O}\in\mathbb{R}^{H\times W\times 3}$ and $F_{S}\in\mathbb{R}^{H\times W\times 3}$ for each filter $t\in\{\text{WST},\text{Canny},\text{Haar},\text{HOG},\text{Grad}\}$, defined as:
$F_{O}=I_{O}+\alpha*\mathbb{FAM}^{t}(I_{O}),$ (1)
$F_{S}=I_{S}+\alpha*\mathbb{FAM}^{t}(I_{S}),$ (2)
$t\in\{\text{WST},\text{Canny},\text{Haar},\text{HOG},\text{Grad}\},$ (3)
IV-B3 Cross-modal Mamba Interaction Module
Existing multi-source fusion methods primarily emphasize spatial-level feature integration, often neglecting the substantial domain discrepancies between heterogeneous modalities such as optical and SAR images. These domain shifts pose a significant challenge to effective cross-modal fusion. The Mamba framework [10], with its state-space modeling and efficient sequence processing capabilities, has demonstrated strong potential for capturing long-range dependencies. This makes it well-suited for mitigating inter-modal inconsistencies. However, existing Mamba-based fusion methods [61, 6] typically involve intricate architectures and computationally intensive interaction designs, which limit their scalability and practical deployment. To overcome these limitations, we propose the Cross-modal Mamba Interaction Module (CMIM), a lightweight yet powerful module tailored to align optical and SAR features by leveraging Mamba’s long-sequence modeling strength. Given paired image features ($F_{Oj},F_{Sj}\in\mathbb{R}^{\frac{H}{2^{j}}\times\frac{W}{2^{j}}\times C_{j}},j\in\{3,4,5\}$), we first flatten them into one-dimensional sequences ($X_{j}\in\mathbb{R}^{1\times n}$, $Y_{j}\in\mathbb{R}^{1\times n}$) in row-major order:
This interleaving ensures that corresponding positional features from both modalities are adjacent, enabling Mamba’s state-space modeling to directly capture semantic and structural alignments, thus enhancing domain coherence. The resulting interleaved sequence is subsequently fed into a directional scanning mechanism (see Fig. 5 [15]), which applies Mamba blocks along both horizontal and vertical directions. This bi-directional sequence modeling allows CMIM to capture long-range dependencies while preserving fine-grained local correspondence, facilitating deeper cross-modal interaction. Finally, the output sequence is reshaped back into the 2D spatial domain and fused with residual skip connections to retain original modality-specific information. Through this design, CMIM effectively bridges modality gaps and enhances the robustness of multi-source object detection under challenging remote sensing conditions.
$Z^{scan}_{j}=\mathbb{SCAN}^{m}(Z^{IIR}_{j}),$ (7)
$m\in\{\text{Bid, Z-Order, Zigzag, Hilbert}\}.$ (8)
@@ IV-B3 Cross-modal Mamba Interaction Module
TABLE II: Quantitative results on the M4-SAR. O and S denote optical and SAR data, respectively. Param. denotes the parameter count of the model. Inf.T indicates inference time per image. The best two results are highlighted in Black and mark.

### v3

@@ 2.3 Data Processing
As shown in Fig. 2 (b), we combined the red, green, and blue channels to generate standard optical images using data processing software (ENVI). The optical images were provided in three spatial resolutions: 10m, 20m, and 60m. To highlight the complementary characteristics of optical and SAR images, we retained only the high-resolution (10m) and low-resolution (60m) variants. For the SAR data, we first applied multi-look averaging to the SLC images in both the azimuth and range directions to reduce speckle effects and improve radiometric resolution. Subsequently, filtering techniques were employed to suppress speckle noise. Using the corresponding DEM data, we geocoded and radiometrically calibrated the SAR images and conducted geometric alignment and correction to improve spatial localization precision. After preprocessing, the optical and SAR images were aligned based on their geographic coordinates. As the overlap between the two modalities was partial, the aligned image pairs were cropped into $\times1,024$ patches, and only those with valid overlapping regions were retained for subsequent annotation. Notably, this initial cropping was intended to facilitate efficient manual annotation. After annotation was completed, we further partitioned the annotated images into non-overlapping $\times512$ patches to meet the demands of different downstream applications.
2.4 Data Annotation and Label Mapping
As the acquisition times of Sentinel-1 [torres2012gmes] and Sentinel-2 [drusch2012sentinel] are not perfectly synchronized, instance-level alignment of dynamic targets (e.g., ships) becomes difficult, which compromises the accuracy of label transfer. To address this issue, we focus on static and semantically meaningful targets, such as bridges, harbors, oil tanks, playgrounds, airports, and wind turbines, and annotate them using oriented bounding boxes. SAR’s unique imaging characteristics cause ambiguous target boundaries, challenging accurate annotation by non-experts. In contrast, optical images (cloudless) provide clearer visual cues that facilitate both target identification and annotation transfer. With precise alignment, annotations derived from optical images can be reliably transferred to SAR images, thereby significantly reducing manual annotation costs.
Given the scale of the dataset, fully manual labeling is both time-consuming and labor-intensive. To address this challenge, we propose a semi-supervised annotation strategy, as illustrated in Fig. 2 (c). The procedure involves the following steps: First, 5% of the entire cloudless optical images are manually annotated, ensuring that each selected image contains all six target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. This subset establishes a high-quality foundation for subsequent annotation. Once manually labeled, a detector [Jocher_Ultralytics_YOLO_2023] is trained using this subset (with the same data used for both training and evaluation during the initial phase). The trained model is then employed to generate pseudo-labels for an additional 5% of the data, which are subsequently refined through manual correction and added to the labeled set. This updated set is used to retrain the detector in the next iteration. The annotation process is repeated iteratively until the full dataset is labeled. During this procedure, images without valid target instances are excluded to enhance annotation efficiency of the M4-SAR dataset.
During the annotation process, we use cloud-free images to ensure annotation accuracy. After annotation is complete, the cloud-free images are replaced with the corresponding cloud-containing images. Based on instance-level alignment, we achieve high-quality label transfer, enabling accurate projection of annotation information from optical images to their corresponding SAR counterparts. As a result, we construct M4-SAR, the large-scale dataset for optical-SAR fusion object detection, which offers a solid foundation for subsequent method evaluation and algorithm development. More dataset details are presented in Appendix 7.
2.5 Data Analysis
As shown in Table 1, our proposed M4-SAR dataset comprises 112,174 image pairs, 981,862 instances, and six representative target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. Each image has a resolution of $\times512$ pixels and contains an average of 8.75 instances, highlighting the dense object distribution typical in remote sensing scenes. As illustrated in Fig. 1 (c), rotated bounding boxes offer significantly higher localization accuracy compared to horizontal bounding boxes. The challenge analysis of category-specific attributes in the M4-SAR dataset is presented as follows:
@@ 2.5 Data Analysis
  • Aspect Ratio Challenge: The aspect ratio of an object plays a critical role in detection accuracy. Fig. 1 (b) illustrates the distribution of aspect ratios across categories. When the aspect ratio approaches 1, targets tend to appear square, potentially leading to recognition ambiguities. Conversely, extremely large aspect ratios increase the model’s sensitivity to object shapes, thus posing additional challenges to accurate detection.
@@
  • Angle Diversity Challenge: All target orientation angles are normalized to $\theta$ ($\theta\in[0,\pi/2)$, consistent with YOLO-OBB [Jocher_Ultralytics_YOLO_2023] rules.), and their distributions are visualized in Fig. 3 (a). Notably, offshore wind turbines tend to exhibit square-like appearances, with orientation angles predominantly distributed to 0 and 90 due to the absence of surrounding reference structures and the top-down perspective of satellite imagery. Although oil tanks are circular, their labeled angles are contextually defined based on surrounding spatial features. This highlights the importance of contextual cues in remote sensing images and reflects the task’s sensitivity.
@@
  • Category Imbalance Challenge: Fig. 3 (b) displays the distribution of instances by category, revealing that bridges (43.64%) constitute the largest proportions in the dataset. This imbalance reflects the actual occurrence frequency of typical targets in coastal regions: static infrastructure such as bridges and oil tanks dominates satellite imagery. The relatively low number of harbor (2.83%) instances is primarily due to their extensive spatial coverage, with a single port typically corresponding to a single instance.
  • Scale Inconsistency Challenge: Fig. 3 (c) illustrates the average pixel area of each category, highlighting the substantial variation in object scales across different categories. Such large discrepancies in object size (i.e., object area) introduce a critical and commonly encountered challenge for robust object detection, as models must simultaneously handle extremely small and large targets within a unified framework.
  • Modal Weak Alignment Challenge: The optical and SAR images are aligned only using geographic coordinates, resulting in coarse rather than pixel-level correspondence. Together with non-simultaneous acquisitions and dynamic scene changes, this leads to notable cross-modal misalignment, as reflected by the low Mean Squared Error (MSE: 0.0662), Structural Similarity Index Measure (SSIM: 0.1256), and Mutual Information (MI: 0.1729) scores shown in Fig. 3 (d). This weak alignment poses a significant challenge for multi-source fusion object detection.
@@
Figure 4: (a) Domain gap between optical and SAR images. (b) Overall framework of the proposed end-to-end optical-SAR fusion object detection (E2E-OSDet). Architectural details of the proposed filter augment module (c), cross-modal Mamba interaction module (d), and area-attention fusion module (e).
3 Evaluation Tools and Methods
3.1 Optical-SAR Fusion Object Detection Tools (MSRODet)
@@
Due to the lack of publicly available methods specifically designed for optical-SAR fusion object detection, conducting a fair performance comparison remains challenging. To address this limitation, we propose a Multi-Source Rotated Detector (MSRODet) built upon existing multi-source detection frameworks (e.g., RGB-infrared, RGB-depth). MSRODet integrates several representative methods, including MHFNet [zhang2026mhfnet], CFT [qingyun2022cross], CLANet [he2023multispectral], CSSA [cao2023multimodal], CMADet [song2024misaligned], ICAFusion [shen2024icafusion], and MMIDet [zeng2024mmi], thereby providing a standardized and accessible benchmark for evaluating optical-SAR fusion object detection approaches. As shown in Fig. 4 (a), empirical analysis reveals that significant domain gaps remain a primary obstacle to effective optical-SAR fusion. These discrepancies primarily arise from fundamental differences in visual characteristics and imaging mechanisms between optical and SAR modalities. Consequently, designing detection frameworks specifically tailored to optical-SAR fusion is essential for enhancing overall detection performance.
3.2 End-to-End Optical-SAR Fusion Object Detection (E2E-OSDet)
@@
Multi-source fusion object detection remains underexplored in the context of optical-SAR fusion, and significant limitations persist in addressing domain discrepancies between optical and SAR data. As shown in Fig. 4 (b), to address these challenges, we propose a novel end-to-end optical-SAR fusion object detection framework, termed E2E-OSDet, which systematically addresses cross-modal challenges across three levels: data input, domain alignment, and feature fusion.
Filter Augment Module (FAM): First, to reduce the domain gap between optical and SAR images, we propose a Filter Augment Module (FAM, Fig. 4 (c)) leveraging classical image filtering operators (e.g., HOG [dalal2005histograms], Canny [canny1986computational], Haar [viola2001rapid], Grad [li2024predicting], and WST [mallat2012group]). This module transforms the low-dimensional and sparse SAR representations into a more discriminative high-dimensional space, effectively mitigating the impact of domain differences.
Cross-modal Mamba Interaction Module (CMIM): Second, most existing methods focus on spatial-level fusion but often neglect the discrepancies in the feature domain. Mamba [gu2024mamba] possesses strong state-space modeling capabilities, which enables effective cross-modal sequence-level interactions and alleviates feature domain bias. However, current Mamba-based methods [zhan2025mambasod, dong2025fusion] are hindered by complex architectures and intricate interaction mechanisms, limiting their practical applicability. To address this, we propose a Cross-modal Mamba Interaction Module (CMIM, Fig. 4 (d)), which enables deep cross-modal interactions through simple and efficient operations.
Area-Attention Fusion Module (AFM): Finally, to further enhance the discriminability of fused features across different spatial regions, we introduce an Area-Attention Fusion Module (AFM). This module captures salient features of local targets and strengthens responses in critical regions, thereby enhancing the effectiveness of feature fusion. The detailed structure of each module is illustrated in Fig. 4 (e). For additional E2E-OSDet details, please refer to Appendix 9.
@@
Table 2: Quantitative results on the M4-SAR. O and S denote optical and SAR data, respectively. Param. denotes the parameter count of the model. Inf.T indicates inference time per image. The best two results are highlighted in Black and mark.
  Method O S Param. FLOPs Inf.T AP50 $\uparrow$ AP75 mAP
(M) $\downarrow$ (G) $\downarrow$ (ms) $\downarrow$ Bri. Har. Oil. Ply. Apr. Win. ALL $\uparrow$ $\uparrow$
R-FCOS [tian2019fcos] $\checkmark$ 31.9 51.6 17.6 32.8 79.8 29.7 67.0 78.9 54.8 57.2 26.4 30.2
R-ATSS [zhang2020bridging] $\checkmark$ 36.0 51.9 19.1 33.2 80.0 29.3 66.6 78.3 54.8 57.0 27.1 30.0
O-RepPoint [li2022oriented] $\checkmark$ 36.6 48.6 16.1 39.7 71.9 21.2 56.3 65.5 53.1 51.3 17.9 26.7
RTMDet [lyu2022rtmdet] $\checkmark$ 8.9 51.0 11.1 34.0 79.8 30.3 61.1 79.1 55.9 56.7 26.5 29.2
PSC [yu2023phase] $\checkmark$ 31.9 51.7 18.3 30.7 79.7 25.9 57.4 78.1 49.4 53.5 26.2 28.5
LSKNet [li2024lsknet] $\checkmark$ 21.8 43.0 20.4 42.9 80.6 34.9 78.4 79.1 54.0 61.6 31.2 34.0
YOLOv5 [Jocher_Ultralytics_YOLO_2023] $\checkmark$ 8.1 12.8 7.2 61.4 88.0 56.6 76.4 91.3 91.8 77.6 59.8 52.9
YOLOv6 [li2022yolov6] $\checkmark$ 16.0 27.7 6.1 57.6 82.1 52.7 67.6 89.4 90.4 73.3 64.9 48.8
YOLOv8 [Jocher_Ultralytics_YOLO_2023] $\checkmark$ 10.0 15.7 6.7 63.6 88.6 57.9 80.0 91.3 92.5 79.0 61.6 54.4
YOLOv9 [wang2024yolov9] $\checkmark$ 6.4 15.1 14.4 61.9 87.7 55.3 74.4 90.0 92.4 76.9 58.9 52.5
YOLOv10 [wang2024yolov10] $\checkmark$ 7.5 14.5 7.8 65.2 89.9 58.8 82.3 90.8 93.5 80.1 62.8 55.7
YOLOv11 [Jocher_Ultralytics_YOLO_2023] $\checkmark$ 9.7 14.4 6.5 64.3 90.0 57.8 80.7 90.9 91.5 79.2 60.5 54.3
YOLOv12 [tian2025yolov12] $\checkmark$ 9.4 14.4 13.4 62.7 88.6 57.1 80.1 91.0 92.7 78.7 61.1 54.0
YOLO26 [Jocher_Ultralytics_YOLO_2023] $\checkmark$ 10.5 15.7 9.1 63.0 88.2 58.3 80.6 90.3 90.1 78.4 62.8 55.4
YOLO-Master [lin2025yolo] $\checkmark$ 29.3 35.6 15.7 64.4 90.0 59.0 80.7 92.3 93.6 80.0 63.9 56.3
R-FCOS [tian2019fcos] $\checkmark$ 31.9 51.6 17.6 23.9 80.1 27.9 60.1 75.8 43.5 51.9 26.4 28.3
R-ATSS [zhang2020bridging] $\checkmark$ 36.0 51.9 19.1 25.3 80.3 26.8 62.3 77.0 43.9 52.6 24.9 27.8
O-RepPoint [li2022oriented] $\checkmark$ 36.6 48.6 16.1 26.2 78.6 21.9 53.7 59.9 48.8 48.2 13.9 20.4
RTMDet [lyu2022rtmdet] $\checkmark$ 8.9 51.0 11.1 16.2 71.0 21.1 39.3 68.9 39.6 42.7 20.9 22.3
PSC [yu2023phase] $\checkmark$ 31.9 51.7 18.3 22.3 80.0 20.6 51.9 75.7 39.4 48.3 24.5 26.5
LSKNet [li2024lsknet] $\checkmark$ 21.8 43.0 20.4 33.4 80.2 28.4 69.3 76.8 51.2 56.6 27.1 30.4
YOLOv5 [Jocher_Ultralytics_YOLO_2023] $\checkmark$ 8.1 12.8 7.2 40.5 85.3 41.0 64.3 85.5 89.9 67.7 44.0 41.3
YOLOv6 [li2022yolov6] $\checkmark$ 16.0 27.7 6.1 36.5 79.4 32.8 53.5 85.7 86.7 62.4 37.6 36.5
YOLOv8 [Jocher_Ultralytics_YOLO_2023] $\checkmark$ 10.0 15.7 6.7 44.9 85.8 42.5 67.9 87.5 89.2 69.6 46.1 43.2
YOLOv9 [wang2024yolov9] $\checkmark$ 6.4 15.1 14.4 43.9 84.8 38.4 61.5 87.9 88.8 67.6 44.2 41.4
YOLOv10 [wang2024yolov10] $\checkmark$ 7.5 14.5 7.8 49.8 87.0 45.5 73.4 87.9 90.4 72.3 50.8 46.2
YOLOv11 [Jocher_Ultralytics_YOLO_2023] $\checkmark$ 9.7 14.4 6.5 46.9 87.2 44.6 71.1 88.8 89.2 71.3 48.6 44.7
YOLOv12 [tian2025yolov12] $\checkmark$ 9.4 14.4 13.0 46.4 87.7 43.6 70.8 90.0 90.7 71.5 49.2 45.4
YOLO26 [Jocher_Ultralytics_YOLO_2023] $\checkmark$ 10.5 15.7 9.1 48.8 88.3 48.9 75.8 88.9 88.9 73.3 53.8 48.2
YOLO-Master [lin2025yolo] $\checkmark$ 29.3 35.6 15.7 49.4 89.3 48.3 75.1 90.6 90.8 73.9 53.1 48.5
COMO [liu2026cross] $\checkmark$ $\checkmark$ 22.6 24.3 37.1 72.3 90.8 60.4 88.7 90.5 96.3 83.2 66.3 57.9
MHFNet [zhang2026mhfnet] $\checkmark$ $\checkmark$ 16.1 26.7 19.2 73.7 91.4 61.8 89.9 90.4 95.6 83.8 65.6 58.0
CFT [qingyun2022cross] $\checkmark$ $\checkmark$ 53.8 31.8 40.6 75.8 92.5 61.3 91.6 90.3 96.3 84.6 68.9 59.9
CLANet [he2023multispectral] $\checkmark$ $\checkmark$ 48.2 53.4 29.1 74.8 92.2 60.7 91.3 91.6 97.2 84.6 68.5 59.6
CSSA [cao2023multimodal] $\checkmark$ $\checkmark$ 13.5 21.4 12.3 73.3 91.7 59.3 88.9 91.6 95.8 83.4 66.4 58.0
CMADet [song2024misaligned] $\checkmark$ $\checkmark$ 41.5 31.8 46.7 70.9 90.7 52.0 86.4 91.7 97.1 81.5 63.5 55.7
ICAFusion [shen2024icafusion] $\checkmark$ $\checkmark$ 29.0 26.7 23.6 74.7 91.9 60.9 91.0 91.8 96.7 84.5 67.3 58.8
MMIDet [zeng2024mmi] $\checkmark$ $\checkmark$ 53.8 31.8 41.9 74.9 92.6 61.1 91.7 91.4 97.0 84.8 68.6 59.8
E2E-OSDet $\checkmark$ $\checkmark$ 24.7 40.7 20.9 77.7 90.7 64.3 91.8 92.1 97.8 85.7 70.3 61.4
4 Experiments and Analysis
4.1 Experiment Setup and Implementation Details